OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#12372 — Optical loop Lille
Incident Report for Web Cloud
Resolved
We have a cut-off in the optics between NRA - ASC59 and TRI59 since 11:55 AM.

NRAs - ASC59, HEL59, LQN59, LZI59 and TRI59 are operating under protection.

A ticket has been opened with the fiber provider.

Update(s):

Date: 2015-01-14 06:31:59 UTC
The incident is completed.Security is operational again.

The operator teams completed the welds on the optical link.

The issue is due to fire incident on the technical room.
Posted Jan 13, 2015 - 11:25 UTC